# Observation of non-Hermitian dislocation bound states and dislocation skin effects

### v1

@@ Abstract
The confluence of Non-Hermitian (NH) topology and crystal defects has culminated significant interest, yet its experimental exploration has been limited due to the challenges involved in design and measurements. Here, we showcase experimental observation of NH dislocation bound states (NHDS) and the dislocation-induced NH skin effect in two-dimensional acoustic NH Chern lattices. By embedding edge dislocations in such acoustic lattices and implementing precision-controlled hopping and onsite gain/loss via active meta-atoms, we reveal robust defect-bound states localized at dislocation cores within the line gap of the complex energy spectrum. These NHDS survive against moderate NH perturbations but gradually delocalize and merge with the bulk (skin) states as the system arrives at the shore of fostering exceptional points in the Brillouin zone under periodic (open) boundary conditions. Furthermore, our experiments demonstrate that the dislocation core can feature weak NH skin effects when its direction is perpendicular to the Burgers vector in periodic systems. Our findings pave an experimental pathway for probing NH topology via lattice defects and open new avenues for defect-engineered topological devices.
preprint: APS/123-QED
@@
Introduction. Crystalline defects, such as dislocations [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14] and disclinations [15, 16, 17, 18, 18, 17], play pivotal roles in the identification of topological crystals in nature and topological metamaterials in laboratory [19, 20, 21, 22]. In Hermitian topological systems, these defects can host robust bound states, protected by a combination of crystalline symmetry and topological invariant as dictated by the bulk-defect correspondence, analogous to the celebrated bulk-boundary correspondence. The emergence of non-Hermitian (NH) physics, where gain/loss and nonreciprocity play significant roles, introduces radical departures from conventional Hermitian behavior [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33]. For instance, NH systems often exhibit phenomena such as the NH skin effect (NHSE), in which a macroscopic number of eigenstates accumulate at the boundaries, obscuring the traditional bulk-boundary signature [34, 35, 36, 37, 38, 39, 40, 41, 42]. These effects necessitate alternative strategies for detecting and characterizing NH topological phases.
@@
Recent theoretical advances have identified crystalline defects as robust local topological probes in NH systems, potentially overcoming the limitations imposed by the NHSE [38, 43]. In particular, Panigrahi et al. theoretically showed that dislocations in a NH Chern insulator can host robust in-gap bound states, protected by pseudo-particle-hole symmetry, whose spatial localization and stability are governed by both NH parameters and the defect geometry [38]. Moreover, an interplay between the dislocation and NHSE can give rise to a dislocation-induced NHSE (D-NHSE), manifesting a typical skin accumulation around the defect core under periodic boundary conditions (PBCs) [36, 37, 38]. However, experimental observation of these fascinating phenomena has remained far from reality due to the requirement for delicate defect engineering combined with precision-controlled NH perturbations, along with the inherent difficulty in probing complex energy spectrum [44, 45, 46, 47, 48, 49].
Figure 1: Schematic implementation of a NH Chern insulator [Eq. (1)] in the presence of an edge dislocation-antidislocation pair with Burgers vectors $\mathbf{b}=\pm\mathbf{e}_{x}$. Each unit cell consists of two sites labeled by A (red ball) and B (blue ball). Solid bars of the same color represent the reciprocal hopping amplitudes of equal strength, and those with arrows indicate nonreciprocal hopping amplitudes. Bottom left inset shows the top view, where circles represent unit cells.
Figure 2: Hermitian ($\mathbf{h}=0$) acoustic Chern insulators in the presence of an edge dislocation-antidislocation pair under periodic boundary conditions (PBCs), showing (a)-(e) experimental observations and (f)-(j) theoretical computations. Energy spectrum [(a) and (f)], amplitude distributions of right eigenstates of the dislocation modes, shown by red dots in (a) and (f) within unit cells [(b) and (g)] and all the states [(c) and (h)] in the ${\rm M}$ phase with $t_{0}=-m_{0}=3\,\mathrm{Hz}$. Energy spectra [(d) and (i)] and amplitude distributions of all the right eigenstates [(e) and (j)] in the $\Gamma$ phase with $t_{0}=m_{0}=3\,\mathrm{Hz}$.
@@
In this work, we circumvent these challenges by embedding edge dislocations into a two-dimensional (2D) acoustic lattice implemented using a coupled acoustic cavity system (CACS) and by introducing controlled NH perturbations via active meta-atoms  [35, 50, 51, 52, 53, 54, 49]. By employing a recently proposed Green’s function approach [49], we directly measure the complex-valued energy spectrum along with both left and right eigenstates of the system. Our experimental platform thereby enables the observation of NH dislocation states (NHDS) emerging within the line gap of the complex energy spectrum, as well as the D-NHSE. We further demonstrate that while the NHDS persist under moderate NH perturbations, they lose support in the system with the appearance of the exceptional points (EPs) therein. As this threshold is approached from the weak NH perturbation side, NHDS gradually delocalize and merge into the bulk (skin) states under periodic (open) boundary conditions; a process coinciding with the closure of the NH band gap, in quantitatively agreement with theoretical predictions. Overall, our work bridges the gap between NH band topology and defect physics, establishing crystal defects as universal tools for probing NH topological matter.
Model and design. We consider a 2D NH Chern insulator whose Bloch Hamiltonian is expressed as [27, 38, 55]
@@
where $\mathbf{d}(\mathbf{k})=\left(t_{0}\sin k_{x},t_{0}\sin k_{y},t_{0}[\cos k_{x}+\cos k_{y}]-m_{0}\right)$, the vector Pauli matrix $\bm{\sigma}=(\sigma_{x},\sigma_{y},\sigma_{z})$, $\mathrm{i}=\sqrt{-1}$, and we set the lattice spacing $a=1$. The Hermitian part is denoted as $H_{\mathrm{H}}(\mathbf{k})=\bm{\sigma}\cdot\mathbf{d}(\mathbf{k})$ and the anti-Hermitian operator is $\mathrm{i}\bm{\sigma}\cdot\mathbf{h}$ with $\mathbf{h}=(h_{x},h_{y},h_{z})\in\mathbb{R}^{3}$. The phase diagram of this model is detailed in the Supplemental Material (SM) [56]. In this model, an edge dislocation-antidislocation pair is introduced by removing two unit cells located at opposite ends along the $y$-direction and subsequently reconnecting the sites along the $x$-direction, see Fig. 1. As a result, any closed path encircling the dislocation core lacks a translation by the Burgers vector $\mathbf{b}=\pm\mathbf{e}_{x}$.
@@
We implement this NH Chern insulator using a CACS. In our design, each acoustic cavity, which supports a dipole resonant mode at the frequency $\omega_{0}=1040\,\mathrm{Hz}-4.5\mathrm{i\,Hz}$, serves as a lattice site. See SM for further details [56]. Here, the imaginary part of $\omega_{0}$ represents the intrinsic background loss. As illustrated in Fig. 1, each unit cell comprises two coupled cavities labeled by A and B. The intra-unit-cell coupling between these two cavities is rendered by the nonreciprocal NH perturbations, yielding asymmetric hopping amplitudes $\mathrm{i}h_{x}-h_{y}$ and $\mathrm{i}h_{x}+h_{y}$, while the inter-unit-cell hopping amplitudes are reciprocal. Both the nonreciprocal and reciprocal couplings are realized via detector-source pairs (i.e., microphone-loudspeaker systems) [49, 54]. For instance, the hopping from site 1 to site 2 is achieved by detecting the acoustic pressure at site 1 and then driving the loudspeaker at site 2 accordingly. In addition, to accurately tune the staggered onsite potentials ±$\pm m_{0}$ and the distributed gain/loss corresponding to ±$\pm\mathrm{i}h_{z}$, we establish a feedback loop within each cavity; a microphone detects the local acoustic pressure and the corresponding loudspeaker modulates the response. The magnitude and phase of each microphone-loudspeaker pair are meticulously adjusted to realize the desired hopping parameters and onsite potentials.
@@
Measurements. Probing the wave dynamics in NH lattices poses unique challenges due to the presence of complex-valued eigenenergies, which manifest as complex poles in the Green’s functions of the system. Conventional single-site excitation techniques, relying on the real or simple complex frequency excitations cannot selectively target these poles [45, 44, 46, 47, 48, 54, 49]. To overcome this limitation, we employ a Green’s function-based measurement method [49]. In our experiments, a source (loudspeaker) is sequentially activated at every site, depicted in Fig. 1, and the resulting frequency responses are simultaneously captured by an array of microphones. This procedure yields the full Green’s function $G(\omega)$ as a function of the excitation frequency $\omega$, from which the complex-valued eigenenergies and corresponding right and left eigenstates are extracted by leveraging the established correspondence with the real-space Hamiltonian.
Figure 3: Experimental observation on an acoustic Chern lattice with an edge dislocation-antidislocation pair under PBCs and moderate NH perturbations for $t_{0}=-m_{0}=3\,\mathrm{Hz}$. (a) Complex energy spectrum for $\mathbf{h}=(1.8\,\mathrm{Hz},0,0)$. (b) Total amplitude of two right eigenvectors of NHDS marked by red dots in (a). (c) Total amplitude of all the right eigenstates from (a). Panels (d)-(f) [(g)-(i)] are analogous to (a)-(c), respectively, but with $\mathbf{h}=(0,1.8\,\mathrm{Hz},0)$ [$\mathbf{h}=(0,0,1.8\,\mathrm{Hz})$].
@@
Parameter space. To ensure a comprehensive observation, we choose four distinct values of each NH perturbation in experiments; $h_{y}=h_{z}=0$ and $h_{x}/t_{0}=0.3,0.6,0.9,1.1$, with analogous values for $h_{y}$ and $h_{z}$. We also include the Hermitian case ($\mathbf{h}=0$), resulting in a total of 13 configurations. For each configuration, experiments were performed under open boundary conditions (OBCs) and PBCs in two distinct phases; the ${\rm M}$ phase with $t_{0}=-m_{0}=3\,\mathrm{Hz}$ and ${\bf K}=(\uppi,\uppi)$, and the $\Gamma$ phase with $t_{0}=m_{0}=3\,\mathrm{Hz}$ and ${\bf K}=(0,0)$. Here, ${\bf K}$ is the band inversion momentum. In total, 52 cases were explored, with all experimentally measured complex-valued energy spectra, left and right eigenstates, and corresponding theoretical predictions detailed in the SM [56].
Observations of NHDS and D-NHSE. We begin our investigation by examining the dislocation bound states in the Hermitian limit ($\mathbf{h}=\mathbf{0}$). Figure 2 displays the experimental results for both the ${\rm M}$ and $\Gamma$ phases. The measured complex energy spectra and corresponding right eigenstates exhibit excellent agreement with theoretical predictions. Namely, only the ${\rm M}$ phase hosts mid-gap dislocation states as ${\bf K}\cdot{\bf b}=\pm\pi$ (nontrivial) around the defect core therein, whereas the $\Gamma$ phase is devoid of such defect modes for which ${\bf K}\cdot{\bf b}=0$ (trivial) [1, 4, 5]. The right eigenstate weight at the $n$-th unit cell is defined as $\Psi_{n}^{\mathrm{R}}=\sum_{m}\Bigl{(}\absolutevalue{\psi_{m,\mathrm{A}}^{\mathrm{R}}}^{2}+\absolutevalue{\psi_{m,\mathrm{B}}^{\mathrm{R}}}^{2}\Bigr{)},$ where the index $m$ sums over either the two states near zero energy corresponding to the dislocation states [Figs. 2(b) and 2(g)] or all states [Figs. 2(c), 2(e), 2(h), and 2(j)]. In this notation, $\psi_{m,\xi}^{\mathrm{R}}$ denotes the right eigenstate for the $m$th state at site $\xi$ within the unit cell, with $\xi={\rm A},{\rm B}$. Notably, PBCs are imposed in Fig. 2 to suppress midgap edge states, thereby enhancing the visibility of dislocation states. SM shows corresponding results under OBCs [56].

### v2

# Observation of dislocation bound states and skin effects in non-Hermitian Chern insulators

@@ Abstract
The confluence of non-Hermitian (NH) topology and crystal defects has culminated significant interest, yet its experimental exploration has been limited due to the challenges involved in design and measurements. Here, we showcase experimental observation of NH dislocation bound states (NHDS) and the dislocation-induced NH skin effect in two-dimensional acoustic NH Chern lattices. By embedding an edge dislocations-antidislocation pair in such acoustic lattices and implementing precision-controlled hopping and onsite gain/loss via active meta-atoms, we reveal robust defect-bound states localized at dislocation cores within the line gap of the complex energy spectrum. We experimentally identify the emergence of bulk exceptional points (EPs) via spectral coalescence and phase rigidity analysis. We demonstrate that the NHDS survive against moderate NH perturbations but gradually delocalize and merge with the bulk (skin) states driven by these EPs under periodic (open) boundary conditions. Furthermore, our experiments demonstrate that the dislocation core can feature weak NH skin effects when its direction is perpendicular to the Burgers vector in periodic systems. Our findings, therefore, pave an experimental pathway for probing NH topology via lattice defects and open new avenues for defect-engineered topological devices.
keywords:
Non-Hermitian physics, topological acoustics, dislocation, defect
Jia-Xin Zhong Bitan Roy* Yun Jing*
{affiliations}
J.-X. Zhong, Y. Jing
Graduate Program in Acoustics, The Pennsylvania State University, University Park, PA 16802, USA
Email Address: yqj5201@psu.edu
B. Roy
Department of Physics, Lehigh University, Bethlehem, Pennsylvania 18015, USA
Email Address: bitan.roy@lehigh.edu
1 Introduction
@@
Crystalline defects, such as dislocations [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14] and disclinations [15, 16, 17, 18], play pivotal roles in the identification of topological crystals in nature and topological metamaterials in laboratory [19, 20, 21, 22]. In Hermitian topological systems, these defects can host robust bound states, protected by a combination of crystalline symmetry and topological invariant as dictated by the bulk-defect correspondence, analogous to the celebrated bulk-boundary correspondence. The emergence of non-Hermitian (NH) physics, where gain/loss and nonreciprocity play significant roles, introduces radical departures from conventional Hermitian behavior [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33]. For instance, NH systems often exhibit phenomena such as the non-Hermitian skin effect (NHSE), in which a macroscopic number of eigenstates accumulate at the boundaries, obscuring the traditional bulk-boundary signature [34, 35, 36, 37, 38, 39, 40, 41, 42]. These effects necessitate alternative strategies for detecting and characterizing NH topological phases.
@@
Recent theoretical advances have identified crystalline defects as robust local topological probes in NH systems, potentially overcoming the limitations imposed by the NHSE [38, 43]. Fundamentally, such defect topology can be classified into point-gap [36, 37] and line-gap [38] phases. While point-gap topology is typically characterized solely by the NHSE, line-gap topology is particularly intriguing as it can host both the dislocation-induced NHSE (D-NHSE) and topological bound states, besides displaying conventional NHSE. In this context, Panigrahi et al. theoretically showed that dislocations in a line-gap NH Chern insulator can host robust in-gap bound states, protected by pseudo-particle-hole symmetry, whose spatial localization and stability are governed by both NH parameters and the defect geometry [38]. Moreover, an interplay between the dislocation and NHSE can give rise to a D-NHSE, manifesting a typical skin accumulation around the defect core under periodic boundary conditions (PBCs) [36, 37, 38]. However, experimental observation of these fascinating phenomena has remained far from reality due to the requirement for delicate defect engineering combined with precision-controlled NH perturbations, along with the inherent difficulty in probing complex energy spectrum [44, 45, 46, 47, 48, 49].
@@
In this work, we circumvent these challenges by embedding a pair of edge dislocation and antidislocation into a two-dimensional (2D) acoustic lattice implemented using a coupled acoustic cavity system (CACS) and by introducing controlled NH perturbations via active meta-atoms [35, 50, 51, 52, 53, 54, 49]. Furthermore, the active meta-atoms are sophistically tuned to induce imaginary hopping terms, thereby breaking the time-reversal symmetry, an essential requirement for realizaing a Chern insulator. This approach offers a significantly greater flexibility compared to conventional flow-based methods [55, 56]. By employing a recently proposed Green’s function approach [49], we directly measure the complex-valued energy spectrum along with both left and right eigenstates of the system. Our experimental platform thereby enables the direct observation of NH dislocation states (NHDS) emerging within the line gap of the complex energy spectrum, as well as the D-NHSE. We further demonstrate that while the NHDS persist under moderate NH perturbations, they lose support in the system with the appearance of the exceptional points (EPs) therein. To validate this mechanism, we explicitly observe the bulk EPs in a pristine lattice, characterized by spectral coalescence and phase rigidity collapse. As this threshold is approached from the weak NH perturbation side, NHDS gradually delocalize and merge into the bulk (skin) states under periodic (open) boundary conditions; a process coinciding with the closure of the NH band gap, in quantitatively agreement with theoretical predictions. Overall, our work bridges the gap between NH band topology and defect physics, establishing crystal defects as universal tools for probing NH topological phases of matter. We emphasize that our realization of a line-gap NH Chern insulator enables the observation of the coexistence of D-NHSE and topological defect-bound states and their interplay with EPs, phenomena that are fundamentally distinct from the D-NHSE characteristic of point-gap systems [57].
2 Results and Discussion
2.1 Model and Design
We consider a 2D NH Chern insulator whose Bloch Hamiltonian is expressed as [27, 38, 58]
@@ 2.1 Model and Design
where $\mathbf{d}(\mathbf{k})=\left(t_{0}\sin(k_{x}a),t_{0}\sin(k_{y}a),t_{0}[\cos(k_{x}a)+\cos(k_{y}a)]-m_{0}\right)$, the vector Pauli matrix $\bm{\sigma}=(\sigma_{x},\sigma_{y},\sigma_{z})$, $\mathrm{i}=\sqrt{-1}$ acts on two sublattices a and b, and we set the lattice spacing $a=1$. The Hermitian part is denoted as $H_{\mathrm{H}}(\mathbf{k})=\bm{\sigma}\cdot\mathbf{d}(\mathbf{k})$ and the anti-Hermitian operator is $\mathrm{i}\bm{\sigma}\cdot\mathbf{h}$ with $\mathbf{h}=(h_{x},h_{y},h_{z})\in\mathbb{R}^{3}$. In this model, an edge dislocation-antidislocation pair is introduced by removing two unit cells located at opposite ends along the $y$-direction and subsequently reconnecting the sites along the $x$-direction, see Fig. 1A. As a result, any closed path encircling the dislocation core lacks a translation by the Burgers vector $\mathbf{b}=\pm\mathbf{e}_{x}$. The phase diagram of this model is shown in Figs. 1B and 1C for nonzero $h_{x}$ or $h_{y}$ and $h_{z}$, respectively. The NH Chern number is C=1C=-1 in the red-shaded region and $C=1$ in the blue-shaded region, while it is trivial outside these areas. See Supplemental Information (SI) Appendix Sec. S1 for details. The system supports topologically protected dislocation states only in the red-shaded regions [38]. In the Hermitian limit ($\mathbf{h}=0$) the above model features two distinct topological phases for $0<m_{0}/t_{0}<2$ and $2<m_{0}/t_{0}<0$, respectively, featuring band inversion at the $\Gamma$ and M points of the Brillouin zone, therefore, named $\Gamma$ phase and M phase.
We implement this NH Chern insulator using a CACS, as illustrated in Fig. 2, which details the experimental setup and parameter tuning methodology. See Experimental section along with SI Appendix Sec. S2 for experimental details, including overall structure and tuning of onsite potentials and hoppings. In our design, each acoustic cavity functions as a lattice site, configured to support a dipole resonant mode at the frequency $\omega_{0}=1040\,\mathrm{Hz}-4.5\mathrm{i}\,\mathrm{Hz}$, where the imaginary component represents the intrinsic background loss. It is important to note that $\omega_{0}$ denotes the reference onsite potential engineered for the NH Chern lattice, rather than the intrinsic resonant frequencies of the as-fabricated cavities. In practice, the natural frequencies of the original cavities deviate significantly due to inevitable fabrication errors and installation imperfections (see SI Appendix Fig. S5). This inherent disorder poses a major challenge to the observation of precise topological phenomena. To overcome this, we employ an active self-feedback mechanism (Fig. 2B) where the signal picked up by a microphone is fed back to a speaker within the same cavity to precisely calibrate and lock the resonance of all 56 cavities to the uniform reference $\omega_{0}$ (see SI Appendix Sec. S2.2 for details). In Fig. 2C1, the target onsite potential is $\omega_{0}-m_{0}+\mathrm{i}\,h_{z}=1043\,\mathrm{Hz}-4.2\,\mathrm{i\,Hz}$, whereas the value extracted by fitting the experimental data is $1042.98\,\mathrm{Hz}-4.21\,\mathrm{i\,Hz}$. In Fig. 2C2, the target onsite potential is $\omega_{0}+m_{0}-\mathrm{i}\,h_{z}=1037\,\mathrm{Hz}-4.8\,\mathrm{i\,Hz}$, with the fitted value of $1037.00\,\mathrm{Hz}-4.81\,\mathrm{i\,Hz}$. Here, $m_{0}=-3\,\mathrm{Hz}$ and $h_{z}=0.3\,\mathrm{Hz}$. These two representative examples demonstrate the high precision of our onsite-potential tuning. Further statistical data regarding the distribution precision and long-term stability of the onsite potentials across all 56 cavities are presented in SI Appendix Sec. S2.5.
@@
As illustrated in Fig. 1A, each unit cell comprises two coupled cavities labeled by ‘a’ and ‘b’. The intra-unit-cell hopping between these two cavities is rendered by the nonreciprocal NH perturbations, yielding asymmetric hopping amplitudes $\mathrm{i}h_{x}-h_{y}$ and $\mathrm{i}h_{x}+h_{y}$, while the inter-unit-cell hopping amplitudes are reciprocal. Both the nonreciprocal and reciprocal hoppings are realized via detector-source pairs (i.e., microphone-loudspeaker systems) [49, 54]. For instance, as shown in Fig. 2D, the hopping from site 1 to site 2 is achieved by detecting the acoustic pressure at site 1 and then driving the loudspeaker at site 2 accordingly. Pure imaginary hopping terms ($\pm\mathrm{i}t_{0}/2$), required to break time-reversal symmetry for realizaing a Chern insulator, are similarly implemented via active components. The magnitude and phase of each microphone-loudspeaker pair are meticulously adjusted to realize the desired hopping parameters and onsite potentials. For configuration in Fig. 2E, the target hopping strength is $\kappa_{0}=t_{0}/2=1.5\,\mathrm{Hz}$, and the retrieved value through fitting experimental results is $\kappa_{0}=1.51\mathrm{\,Hz}+0.01\mathrm{i\,Hz}$. For configuration Fig. 2F, the target hopping strength is $\kappa_{0}=-h_{y}+\mathrm{i}h_{x}=-0.3\,\mathrm{Hz}+0.3\mathrm{i\,Hz}$, and the retrieved value through fitting experimental results is $\kappa_{0}=-0.30\mathrm{\,Hz}+0.30\mathrm{i\,Hz}$. These two representative examples demonstrate the high precision of our hopping strength tuning. Further statistical data regarding the distribution precision and long-term stability of the hopping strengths across all 56 cavities are presented in SI Appendix Sec. S2.5.
We highlight that this fully programmable active coupling scheme offers distinct capabilities compared to passive acoustic interconnects or topolectrical circuit. Unlike passive tubes typically restricted to reciprocal Hermitian couplings, our active meta-atoms enable the independent tuning of nonreciprocal terms as well as complex-valued hoppings that explicitly break time-reversal symmetry, which are essential for accessing the line-gap Chern phase. Furthermore, regarding electric circuits, we note a fundamental distinction in physical realization. While circuits are powerful tools for emulating the Hamiltonian matrix structure [59], our platform realizes the Hamiltonian governing the propagation of an actual physical wave field. Consequently, our observation of defect-core localization and NHSE serves as a direct demonstration of these phenomena in a wave medium, offering valuable insights transferable to other wave-based platforms such as photonics and mechanics.
2.2 Experimental Measurements
@@
Probing the wave dynamics in NH lattices poses unique challenges due to the presence of complex-valued eigenenergies, which manifest as complex poles in the Green’s functions of the system. Therefore, conventional single-site excitation techniques, relying on the real or simple complex frequency excitations, cannot selectively target these poles [45, 44, 46, 47, 48, 54, 49]. To overcome this limitation, we employ a Green’s function-based measurement method [49]. In our experiments, a source (loudspeaker) is sequentially activated at every site, depicted in Fig. 1A, and the resulting frequency responses are simultaneously captured by an array of microphones. This procedure yields the full Green’s function $G(\omega)$ as a function of the excitation frequency $\omega$. Representative raw spectral responses measured at distinct lattice sites are provided in SI Appendix Sec. S3. We note that directly identifying the topological features (NHDS or D-NHSE) from these raw point-to-point spectra is challenging due to the intrinsic background loss, which obscures the specific complex poles. However, by analyzing the full measured Green’s function matrix, the complex-valued eigenenergies and corresponding right and left eigenstates are accurately extracted. Importantly, this Green’s-function-based approach is universally applicable to general linear wave systems, including those utilizing passive coupling mechanisms (e.g., acoustic tubes). To demonstrate this generality, we successfully applied this method to extract the complex spectrum and eigenstates of a minimal two-cavity system connected by passive tubes (see detailed validation in SI Appendix Sec. S2.8).
2.3 Parameter space
@@
To ensure a comprehensive observation, we choose four distinct values of each NH perturbation in experiments; $h_{y}=h_{z}=0$ and $h_{x}/t_{0}=0.3,0.6,0.9,1.1$, with analogous values for $h_{y}$ and $h_{z}$. We also include the Hermitian case ($\mathbf{h}=0$), resulting in a total of 13 configurations. For each configuration, experiments were performed under open boundary conditions (OBCs) and PBCs in two distinct phases; the $\rm M$} phase with $t_{0}=-m_{0}=3\,\mathrm{Hz}$ and ${\bf K}=(\uppi,\uppi)$, and the $\Gamma$ phase with $t_{0}=m_{0}=3\,\mathrm{Hz}$ and ${\bf K}=(0,0)$. Here, $\bf K$} is the band inversion momentum. In total, 52 cases were explored, with all experimentally measured complex-valued energy spectra, left and right eigenstates, and corresponding theoretical predictions detailed in the SI Appendix Sec. S6 and Figs. S20–S71.
2.4 Observations of bulk EPs
Before investigating the interplay between NH topology and dislocations, it is crucial to experimentally validate the emergence of EPs in the bulk lattice. We characterize the spectral properties of a pristine $4\times 4$ lattice without dislocations under PBCs. Figure 3 presents the experimental observation of EP signatures as the NH perturbations ($h_{x},h_{y},h_{z}$) are tuned across the theoretical critical value ($h_{j}/t_{0}=1,j=x,y,z$). Figures 3A-F display the evolution of the complex energy spectra. The discrepancies between the experimental data and the theoretical predictions may arise from disorder in the onsite potentials and hopping amplitudes induced by experimental imperfections (see SI Appendix Sec. S4 for details). As the system approaches $h_{\mathrm{c}}$, we observe a distinct coalescence of eigenenergies in the complex plane, signaling the proximity to EPs.
To quantitatively verify the EP characteristics, we introduce the phase rigidity $r_{n}$ of the $n$th eigenstate, defined as
$r_{n}=\frac{|\langle\psi_{n}^{\mathrm{L}}|\psi_{n}^{\mathrm{R}}\rangle|}{\sqrt{\langle\psi_{n}^{\mathrm{L}}|\psi_{n}^{\mathrm{L}}\rangle\langle\psi_{n}^{\mathrm{R}}|\psi_{n}^{\mathrm{R}}\rangle}},$ (2)
[…]
Furthermore, Figures 3H1-H3 reveal that the real-energy line gap ($\Delta E$, defined by the separation of two dashed lines in Fig. 3A1) starts to close at this critical point. This bulk gap closure provides the physical mechanism for the “melting” of the NHDS observed in the subsequent sections. Since the topological protection of NHDS relies on the existence of an open line gap, its closure at the EP naturally leads to the delocalization of defect modes and their absorption into the bulk continuum.
2.5 Observations of NHDS and D-NHSE
[…]